RAZR Industries issued a 25 year bond, with a coupon of 8%, making semi-annual payments and a par value of $1,000. The bond currently sells for 108% of par. The company's tax rate is 21%.

1) Calculate the before tax cost of debt.

2) Calculate the after tax cost of debt.
Answer #1

Information provided:

Par value= future value= $1,000

Current price= present value= 108%*1,000= $1,080

Time= 25 years*2= 50 semi-annual periods

Coupon rate= 8%/2= 4%

Coupon payment= 0.04*1,000= $40

1.The before tax cost of debt is calculated by computing the yield to maturity.

Enter the below in a financial calculator to compute the yield to maturity:

FV= 1,000

PV= -1,080

PMT= 40

N= 50

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 2.2885.

Therefore, the before tax cost of debt is 2.2885%*2= 4.5770%. 4.58%.

2.After tax cost of debt= before tax cost of debt*(1 – 0.2)

                                            = 0.0458*(1 – 0.21)

                                            = 0.0362*100

                                            = 3.62%.

In case of any query, kindly comment on the solution.